OpenStudy (tlynne16):

Bob has 20 feet of fencing to enclose a rectangular garden. If one side of the garden is x feet long, he wants the other side to be (10 – x) feet wide. What value of x will give the largest area, in square feet, for the garden? A. 3 B. 4 C. 5 D. 6

OpenStudy (phemenway0001):

let A(x) = x(10 - x) ← area = length × width = 10x - x² = -(x² - 10x) then put into vertex form: = -(x² - 10x + (-5)²) + 25 ← completing the square = -(x - 5)² + 25 which is an open downwards parabola with vertex: (5, 25) so answer: C. 5 and max area is 25 ft²